Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 8629.90 nb-1.

Monday June 30
The day shift began with Operations monitoring store 2736, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 1:05 PM, a Main Injector (MI) expert began conducting a study that required MiniBooNE beam to be off. Operations returned beam to MiniBooNE at 1:42 PM.

A Linac expert made an adjustment to some equipment and then tuned the Linac beam at 4:48 PM.

A Tevatron (TeV) lead failure at sector F4 caused the loss of store 2736 at 6:36 PM. A cryo system valve (EVDH) opened and then closed on its own accord. Operations notified cryo experts.

Operations put the TeV through a dry squeeze at 7:36 PM.

Booster suffered a vacuum burst at 8:02 PM.

Since the colliding beam is gone, cryo technicians will change a belt on the TeV sector D4 wet engine. They started work at 8:11 PM and finished an hour later.

Operations established store 2737 at 10:24 PM. Its initial luminosity was 26.56E30. Pbar resumed stacking.

Tuesday July 1
The midnight shift began with Operations monitoring store 2737, with Pbar stacking, and with MiniBooNE taking beam.

A Linac quadrupole (L:Q307) power supply tripped off at 12:09 AM and again five minutes later. Operators changed a control card.

L:Q307 tripped off at 12:38 AM. It reset okay.

At 3:36 PM, a Booster kicker (MKS05) tripped off and operators could not get it working again. They timed in another kicker (MKS04) at 4:03 AM.

Operators had to reset the breaker for Q307 at 4:31 AM and then again at 4:49 AM.

Q307 tripped off at 6:31 AM. Operators replaced all of the control cards.

Two Booster RF stations (BRF5 & 6) tripped off at 8:39 AM due to a vacuum pump tripping off. Everything reset okay.

A Booster expert began investigating possible damper problems at 9:07 AM.

A Booster dipole magnet power supply tripped off due to over temperature at 9:33 AM. Operators placed a temporary fan on the supply and had it operating by 9:50 AM, but the supply continued to run warm.

A Booster expert discovered and repaired a bad cable on kicker MKS05 at 10:16 AM. The expert checking on the Booster dampers said that they were operating normally.

A Switchyard expert began a study period at 3 PM.

Booster experts began checking all the Booster kickers at 4 PM.

Operations turned off MI beam and began preparing it for access at 4:04 PM to repair the PA for a RF station (MIRF12).

Operations turned off the H- Source at 4:06 PM so experts could perform maintenance. They were finished by 4:30 PM.

Vacuum techs changed out two Booster vacuum pumps at 4:30 PM.

Operations terminated store 2737 at 5:13 PM.

The D-zero (D0) experiment accessed their collision hall at 5:37 PM. They were out of access by 6 PM.

MI was ready for beam by 6:45 PM. Unfortunately there was a problem with the motor generator in the H- Source. Operators accessed the dome, but couldn't find any obvious problems. An expert was called in and arrived at 8:20 PM. Another expert called saying that a gas valve had failed and he was coming in.

Operators began preparing the I- Source for operation at 9 PM. Another expert was called in. Experts began adjusting the I- Source at 11:48 PM.

Wednesday July 2
The midnight shift began with the I- Source operating, but still warming up (which means a lower intensity of beam. Operators are just starting to resume sending beam to MiniBooNE.

Pbar resumed stacking at a reduced rate at 12:09 AM.

A MI expert arrived at 1:15 AM to begin coalescing studies. She finished her study at 2:02 AM.

A Booster RF station (BRF9) tripped off at 2:20 AM. Operators had it back up by 2:38 AM, but it was operating poorly.

At 2:34 AM, Operations reported struggling with MiniBooNE beam for the last forty minutes. Operators continued to make adjustments.

Operations put the TeV through a dry squeeze at 2:39 AM.

Operations put the TeV into shot set up at 3:17 AM.

At 3:31 AM, Operations noted Linac beam problems related to the new source. Operators continued tuning.

During shot setup, at 4:56 AM, a problem occurred with the Pbar sequencer after the fifth transfer of antiprotons. Operators contacted experts and then tried to reinitiate the sequencer. This didn't work as it was supposed to. The twenty-five minute delay of shot setup allowed the intensity of the beam already in the TeV to degrade.

Operations established store 2740 at 5:45 AM. Its initial luminosity was 22.0E30.

Today's Plan
The plan for today is to stack and store. The Recycler access to the Main Injector that was scheduled for today has been canceled.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/2/2003   email Fermilab